# Low Resource Reconstruction Attacks Through Benign Prompts

### v2

@@ 4.3 Comparison with previous attacks
As discussed, to allow comparison with previous attacks we also extracted a list of collocations, similar to the one we extracted from the generic websites we chose, by harvesting e-commerce sites that appeared in previous works. For example, we could extract the collocation car seat cover and compare our approach to previous approach that mines captions from the training set. Fig. 7 provide examples of images and the associated prompts used in each attack.
Webster Ours 1 10 1 10 Refer to caption Refer to caption Refer to caption Refer to caption Tribal Aztec Indians pattern Universal Fit Car Seat Covers Floral Car Seat Covers Refer to caption Refer to caption Refer to caption Refer to caption Skull Floor Pillow Size: 36" Floral Floor Pillow Refer to caption Refer to caption Refer to caption Refer to caption 3D Black & White Skull King Design Luggage Covers 007 Floral Luggage Covers
Figure 7: Comparison of images generated after 1 step and 10 steps across prompts from Webster’s attack and ours. Images from top to bottom: (a) Identified by Webster’s white-box attack but not the black-box attack. (b) Identified by the black-box attack as non-verbatim. (c) Identified as template verbatim.
@@
5 Observed Phenomena in Template-Memorized Images
By reconstructing many template-memorized images, we were able to observe the following phenomena, which, to the best of our knowledge, were not previously discussed in literature.
5.1 Interpolation
An important type of image recreation we observed involves interpolation of several real-world images. Several images that were generated by our process were not necessarily verbatim copy of an existing image, nor did they involve naive copying of the background or foreground. Instead, we could find that certain elements were copied from different sources in a single image. This phenomena is demonstrated in Fig. 9. Such type of copying can be challenging to detect. For example, one of the images that can be seen in Fig. 9 copies a tattoo that appeared in a real-world image, but not the whole image. That means that the copied image and the copying image might not be close at all at any metric, and yet it is clear that certain elements in the picture were blatantly copied. This also raises important questions about the level of granularity at which copying should be defined and evaluated. We speculate that some generated images may be interpolated from numerous source images, making them difficult, or even infeasible, to trace with current methods and resources.
Figure 9: Examples of interpolations observed, Left: "X Essential T-Shirt", right: "X iPhone Case & Cover" The elements in the right columns were also identified in the attack of Somepalli et al. [32]
5.2 Perturbations
We’ve noticed clusters of images that were nearly identical in the template sense but with one or more objects perturbed between semantically similar objects in a fixed location such as lamps and chairs. This is exemplified in Fig. 10. Such perturbations preserve a high CLIP similarity (outside the editable region) between the images. Perturbations somewhat affect the pixel-wise $\ell_{2}$ similarity since they change the pixel values, but still maintain a higher-than-random similarity. While difficult to pinpoint the source of this phenomena, it stands in concurrence with the expected consequences of minor perturbations in the noise latent space, as demonstrated in [23]. It might also stem from the one-step denoising behavior described by [36], as a clear layout of the image after the first step, leaves room only for minor changes in the next steps.
5.3 Leakage
Another interesting phenomenon that we identified is a certain template leakage where an image template belonging in one template set appears is generated by a prompt associated with another template group, or not associated with a template at all. In the image generated from template leakage, the object is roughly overlapping with the edited region, in a visually sensible way. Such an example of suspected leakage is demonstrated in Fig. 11(b) where a template associated with the category "T-Shirt" appeared under "Tank Top". The edited area is a tank top, but the background appeared in the source and other generated images with a T-Shirt.
Inspecting e-commerce websites, though, some image templates are re-used for different product. Thus, some cases of suspected leakage may not be such. For example see the common background between "Canvas Wall Art Print" and "Wall Tapestry" Fig. 11(a). The image template appeared with both categories in the training data.
(a) Background shared between template groups, not suspected for leakage. Left: "Galaxy Wall Tapestry", right: "Galaxy Canvas Wall Art Print".
(b) Suspected leakage from "T-Shirt" to "Tank Top". Left: "Abstract Art Essential T-Shirt", right: "Abstract Art Tank Top".
Figure 11: Examples of templates shared between image categories in the training data, including suspected leakage.
Acknowledgments
Appendix A Synthetic Results
We wanted to verify our hypothesis on how the coupling between text templates and image templates causes partial memorization, therefore we conducted synthetic experiments mimicking the sort of coupling in order to intentionally create memorization. Another phenomenon that we wanted to identify and understand is what we term leakage. We have observed on real-world data that certain images from one category included templates from another category. Unfortunately, we could not verify if this is due to leakage from categories or whether certain templates simply appeared on the training data on different category (a phenomenon that we could identify that sometimes happen). Therefore, through synthetic experiments we validate that indeed leakage may be happening.
For our synthetic experiments, we collected 3 images of a coffee Mug with an iPhone SE where we placed the same coffee mug in 3 different locations in our lab. We then manually created a mask for each image with Photopea and then simply replaced the mask area with a pattern using OpenCV. In this stage, the patterns were crudely overlayed, creating an unnatural appearance that was easy for the model to memorize. We observed template verbatim extraction along with interpolation, perturbations and leakage.
As a second stage, we used mockups with more realistic rendering of the overlaid pattern based on the technique used for actual Print on Demand websites. The mockups were taken from Freepik, to one of the mockups we added in Photoshop 2 elements also taken from Freepik: a pair of leaves and a slice of lemon. We selected 3 mockups and replaced the smart object contents in Photoshop by an automation script with some manual fixes to ensure that the overlay is uniform between the images. Beside the different templates, the rest of the experiment remained the same as the first stage. We did not observed template verbatim extraction, but we did observe interpolation, perturbations, and object memorization.
To validate the coupling between text and images, we divide this experiment into two stages. In each stage of this experiment we created a dataset mimicking the data from design placement websites on images created by us for that purpose such that they didn’t appear in the training set. We selected the category of "Coffee Mug" which did not involve identified templates. Next, we select 3 images of a blank coffee mug and overlay patterns on them as detailed below. The overlaid patterns are 18 patterns with CC0 license downloaded from Freepik [7]. The caption for each image is composed of a short description for the pattern e.g. "Floral" and the category "Coffee Mug". The dataset of 54 image-caption pairs were used to fine-tune Stable Diffusion by LoRa [22]. We then generated images from the trained model using the prompt template “skg Coffee Mug” where “skg” had been replaced with a pattern description, including patterns that were not in the training set.
In a addition to generating the collocation "Coffee Mug", we generate images from our fine tuned model through semantically similar and semantically dissimilar to “Coffee Mug”: “Tea Cup” and “T-Shirt”, Accordingly. Through these additional prompts we could identify that leakage indeed happens. Specifically for the prompts "skg Tea Cup", the backgrounds of all of the images are templates interpolation or at least similar colors/textures, the tea cups are also in top view which was not seen during training and did not appear in the coffee mug generated images (only the 3 views from the train set), and also in different shapes.
The results of our synthetic experiments support our assumption on how the template text-image structure in the training set encourages template-memorization, as well as demonstrating the observed phenomena we discussed at 5.
@@ Appendix A Synthetic Results
Appendix B Traces of memorization in Stable Diffusion 3.5
Beside clear template extraction up to perturbations, we also noticed that even in cases where our original attack failed, traces of the attacks could still be found and perhaps utilized in future, more refined, attacks. For example, although the beach towel in Fig. 3 could no longer be reconstructed directly, identifying its source via Google Lens revealed that the caption still produces distorted versions of the original image. This indicates that training images from e-commerce websites may continue to pose risks in future iterations of diffusion models.
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 13: Traces of our attack in SD 3.5 (Medium). The first three images were generated using the prompt "BOYOUTH Round Beach Towel,X Beach Mat with Tassels Ultra Soft Super Water Absorbent Multi-Purpose Towel,59 inch-Diameter" Taken from the source image’s product description in Amazon, where X was: "Abstract Art", "Floral", "Galaxy". The image to the right of the vertical line is the source image, which was reconstructed in previous models (see Fig. 3)

### v3

@@ 4.2 Comparison with previous attacks
As discussed, to allow comparison with previous attacks we also extracted a list of collocations, similar to the one we extracted from the generic websites we chose, by harvesting e-commerce sites that appeared in previous works. For example, we could extract the collocation car seat cover and compare our approach to the previous approach that extracts captions from the training set. Figure˜7 provide examples of images and the associated prompts used in each attack.
@@
Demonstration of the one-step denoising behavior presented at [37]can be seen at 8.
Webster Ours
1 10 1 10
Refer to caption Refer to caption Refer to caption Refer to caption
Tribal Aztec Indians pattern Floral Car Seat Covers
Refer to caption Refer to caption Refer to caption Refer to caption
Skull Floor Pillow Floral Floor Pillow
Refer to caption Refer to caption Refer to caption Refer to caption
3D Skull King Luggage Cover Floral Luggage Cover
Figure 8: Comparison between images generated after 1 and 10 steps for Webster’s attack and ours. From top to bottom: (a) identified by Webster’s white-box only, (b) identified as non-verbatim, (c) identified as template verbatim.
4.3 Attacks on State-Of-The-Art and other models
Stable Diffusion v3.5 Flux-Schnell v1.0 MidJourney v6.1
Source Generated Source Generated Source Generated
While we focused our efforts and designed our attack on Stable Diffusion 1.4 (SD 1.4), we could also extract template-memorized images from other popular generative models. Primarily, by using the same collocations identified in SD 1.4. In Figure˜10 we exhibit further images extracted from DeepFloyd IF-XL-I-v1.0 [7] and Midjourney V4 [25]. Both models were released around the time of SD 1.4, and in Figure˜9 we demonstrate further extractions from state-of-the-art (SOTA) models.
  • DeepFloyd IF-XL-I-v1.0: although the generated images exhibited lower contrast, which made clique search impractical, we successfully identified template-memorized images through visual inspection. In addition, we discovered new templates, absent from SD 1.4, using Google Lens searches on manually flagged generations.
    For each prompt and seed, we collected the four default outputs returned by Midjourney, specifying version 4 with the argument --v 4. Despite limited prompt and seed control, we successfully extracted template-like generations from Midjourney. Our results included both direct template replications from SD 1.4 and interpolated reconstructions (seeSection˜5.1)).
    DeepFloyd
    Example results from our attack on DeepFloyd. For each pair, the source image (left) and the generated image (right) are shown.
    Midjourney v4
    Figure 10: Reconstructed images from models of the same generation as SD 1.4. For each pair, the source (left) and generated (right) images are shown.
    For Stable diffusion and flux Schnell we used the Hugging Face checkpoint in the same manner we’ve conducted our main experiments on SD v1.4, while for Midjourney we used the Discord interface as was detailed previously. We successfully demonstrated some memorization, as seen in Figure˜9, though to a lesser extent than in older models and versions.
    SD 3.5 introduced partial mitigation techniques targeting image-text coupling vulnerabilities; however, these measures were not specifically designed to address the template-style coupling exploited by our attack, which relies on concise prompts derived from external sources rather than training data. Our results indicate that SD 3.5 exhibits increased resilience but remains partially vulnerable to template memorization. Representative examples are provided in Figure˜9
    MidJourney V6.1, as discussed, poses additional challenges. However, we applied the same methodology used for V4. While the generated images were typically perturbed (see (Section˜5.2), targeting previously vulnerable benign prompts still yielded images with notable resemblance to the source. For instance, as shown in Figure˜9, the prompt “Abstract Art Essential T-Shirt" (which previously produced near-verbatim copies) continues to generate closely resembling images.
4.4 User Study
To validate our manual annotations of copied images, we conducted a user study, to confirm that the copying is visually apparent to a human observer. Prior work focused on verbatim copies, where metrics like SSCD or edge detection were incorporated in the pipeline. In contrast, our work targets template-level memorization, where the reproduced content may be in the background, or differ in minor structure. This makes the problem of copying ill-posed, and we observed that edge consistency frequently failed to flag images that, to a human observer, were clearly copied. We therefore validated our observations through a user study that confirmed the success of our attack. Participants were shown generated images and potential source images, and were asked to identify copied images. The selected examples included both direct instances of TMI and cases of interpolations and perturbations, as described in Section˜5.1 and Section˜5.2, respectively. Overall, we collected a dataset of 56 source images and corresponding generated counterparts. 70 users answered our questionnaire and each user had to answer 18 to 19 questions. For each generated image, we also selected a control image which was another output produced using the same prompt but a different, randomly chosen seed. We also showed examples where both generated images and none contained copied elements from the source. In each question, participants were shown a triplet consisting of a source image and two generated images, and were asked:
“For each image pair, select which generated image appears to have elements copied from the source.”
Figure 11 presents a summary of the study, showing that user judgments validate our identification of copied content.
The available choices were: A, B, Both, and None. The questionnaires were administered through Google Forms, and the participants non-expert volunteers recruited through personal networks. To verify the diversity of the sample, we also collected age and gender information.
Figure 11: User study results aggregated by our ground-truth labels with cross-participant CI.
4.5 Real-World Benign Prompts
We previously characterized our attack as reflecting the behavior of a naive user who interacts with the model exactly as intended. To further support this claim, we analyze two large-scale datasets of real user prompts [24, 8], collected directly from everyday model usage.
First, in each dataset, we search for prompts containing our identified collocations to validate that indeed such collocations appear in every-day use. In the Midjourney dataset [24], we managed to find 12,706 total prompts (3,883 unique) containing these collocations out of 7.13M samples. In the Stable Diffusion dataset [8], we could find 818 total prompts (660 unique) out of 1.8M samples.
The original images generated from these prompts were unavailable due to link expiration (Midjourney) and dataset scale (DiffusionDB) so we could not estimate how many of these prompts actually yielded TMIs in real-world use cases. Nevertheless, as a preliminary assessment, we conducted a small, low-resource experiment focusing on a limited set of selected candidate prompts From the 3,883 unique Midjourney prompts and 660 unique DiffusionDB prompts that contain our collocations, we select only the shortest prompt, yielding curated subsets of 35 candidate prompts. We then generated images using our previously described pipeline and manually inspect them for the templates we have identified (see A). Despite the modest scale of this experiment, the results were notable: 8 out of 20 Midjourney prompts and 7 out of 15 DiffusionDB prompts produced TMIs. An example prompt and its resulting TMI appear in Figure˜1. Although preliminary, these findings suggest that TMIs may be more common in the wild than previously assumed. A more extensive study is needed to determine their true prevalence and evaluate the seriousness of this phenomenon.
4.6 Comparison with previous works
As discussed, to allow comparison with previous attacks we also extracted a list of collocations, similar to the one we extracted from the generic websites we chose, by harvesting e-commerce sites that appeared in previous works. For example, we could extract the collocation car seat cover and compare our approach to previous approach that mines captions from the training set. Figure˜12 provide examples of images and the associated prompts used in each attack.
Generated Image Source Image Prompt and Source Caption
Caption: Wild Hearts Can’t Be Broken Car Seat Covers For Horse Lovers 170804 – YourCarButBetter.
Caption: Designart “Wide Pathway in Yellow Fall Forest” Landscape Photo Round Metal Wall Art.
4.7 Traces of Memorization in Stable Diffusion 3.5
Beyond the clear instances of template extraction—sometimes preserved only up to perturbations, as shown in Figure˜9—we also found that Stable Diffusion 3.5 exhibits subtler forms of memorization. Even when our original attack no longer yielded a recognizable reconstruction, meaningful traces of the underlying training content persisted, suggesting that future, more targeted attacks may still uncover sensitive material.
These findings highlight the significance of the phenomena described in this work—interpolations, perturbations, and template leakage. Without understanding these behaviors, one might incorrectly assume that failed reconstructions imply the absence of memorization. Instead, these phenomena reveal that memorized content can manifest in isolated visual elements, partially altered fragments, or blended structures rather than as a direct or verbatim template. In other words, the form in which memorization appears can change—even when its presence does not.
A concrete example appears in Figure˜13. Although the image template had no longer been fully reconstructed in Stable Diffusion 3.5, its generation still exhibited recognizable elements, such as a person in a specific pose. Using the template groups we had previously identified for Stable Diffusion 1.4, we were able to match this distorted output back to the same underlying template and corresponding source image. This shows that, even when verbatim extraction is no longer possible, remnants of training images, particularly from large-scale e-commerce data,can survive as structured traces tied to known templates, and thus may continue to pose a risk in future model versions.
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 13: Traces of our attack in SD 3.5 (Medium). The first three images were generated using the prompt "BOYOUTH Round Beach Towel,X Beach Mat with Tassels Ultra Soft Super Water Absorbent Multi-Purpose Towel,59 inch-Diameter" Taken from the source image’s product description in Amazon, where X was: "Abstract Art", "Floral", "Galaxy". The image to the right of the vertical line is the source image, which was reconstructed in previous models (see Figure˜5)
4.8 Synthetic Experiments
To better understand how the coupling between text templates and image templates contributes to partial memorization, we designed a series of synthetic experiments that deliberately recreate this form of coupling. Our goal was twofold: (1) to intentionally induce memorization under controlled conditions, and (2) to investigate another phenomenon we call template leakage. In real-world data, we frequently encountered images where a template from one category appeared under a different category. However, because e-commerce platforms often reuse templates across product types, it was impossible to determine whether these cases reflected true leakage or simply the natural reuse of templates in the training data. Synthetic experiments offered a way to isolate these effects.
Stage 1: Controlled overlays with crude template replacement.
For the first stage, we collected three photographs of a coffee mug next to an iPhone SE, placing the same mug in three different locations in our lab. We then created manual masks for each image using Photopea and replaced the masked regions with simple patterns using OpenCV. At this stage, the overlays were intentionally crude—visually unrealistic and easy for the model to memorize.
Under these conditions, we observed verbatim template extraction, as expected, along with clear instances of interpolation, perturbations, and leakage. The model reproduced not only the underlying structure of the templates but also combined and varied them in ways that matched the phenomena we identified in real-world data.
Stage 2: Realistic mockups mimicking Print-on-Demand rendering.
In the second stage, we constructed a more realistic setup based on Print-on-Demand (PoD) workflows. We selected three mockups from Freepik and replaced their smart-object contents in Photoshop using an automation script, with light manual adjustments to maintain uniformity across images. For one mockup, we additionally inserted two decorative elements—a pair of leaves and a lemon slice—also from Freepik, to mimic the kind of compositional variability seen in real product designs.
With these more naturalistic overlays, we no longer observed verbatim template extraction. However, interpolation, perturbations, and object memorization remained evident, confirming that these behaviors persist even when the templates are more realistically embedded and less trivially memorized.
Testing for template leakage under semantic shifts.
To probe template leakage directly, we generated images not only for the collocation “Coffee Mug” but also for prompts that were semantically similar (“Tea Cup”) and semantically dissimilar (“T-Shirt”). These generations allowed us to separate genuine leakage from simple template reuse.
The results provided clear evidence of genuine leakage. Under prompts such as “skg Tea Cup,” the backgrounds consistently showed interpolations—or, at minimum, strong color and texture echoes—of the coffee-mug templates used during fine-tuning. Even more strikingly, the generated tea cups appeared in top view, a viewpoint that never appeared in the training data and was absent from all coffee-mug generations, which only reproduced the three canonical views present in the training set. The cups also manifested in shapes not seen during training.
Together, these findings demonstrate not only that template leakage occurs, but also that synthetic experiments enable us to disentangle leakage from benign template reuse- revealing how template structures can transfer across prompts and categories under controlled conditions.
@@ 4.8 Synthetic Experiments
5 Observed Phenomena in Template Memorized Images
Building on the successful reconstruction of a large set of template-memorized images, we observed three notable phenomena—Interpolation, Perturbations, and Leakage—which, to the best of our knowledge, have not been previously reported in the literature.
Before moving to describe the phenomena, we first note that the generation of template-memorized content follows a many-to-many structure which we name template groups: a set of collocations are associated with a set of image templates, rather than the one-to-one text-image association typically assumed. This structure was hinted at [37] but not named and studied. We further note that the collocations in the group are either semantically similar (e.g. duvet cover, comforter, throw blanket) or appeared together in product description, as they were listed together as options of the same product (e.g. picnic blanket, beach towel). The template groups we identified are listed at A.
5.1 Interpolation
The first phenomena we observe involves interpolated reconstructions, where generated images blend elements from multiple training examples rather than copying any single image outright. Such partial copying evades standard similarity metrics, as neither the background nor foreground is directly duplicated, but the elements can be easily identified as matching each source by a human observer, as further validation in the user study presented in 4.4. In Figure˜15, for instance, one image contains a tattoo reproduced from a real-world photo, while the hair reproduced from another. The combination of elements from multiple sources raises challenging questions regarding what constitutes a memorized output.
This form of interpolation reveals a deeper behavior: diffusion models appear to internalize, reorganize, and recombine fine-grained visual traits drawn from multiple training images. In our experiments, we were able to identify interpolated reconstructions that clearly fused elements from two—sometimes three—distinct memorized sources.
We further speculate that some generations may blend far more subtle fragments from a broader set of images, producing composites whose lineage spans numerous sources. If such multi-source blending indeed occurs, it becomes increasingly difficult—sometimes effectively impossible—to determine which training images influenced a given output. This, in turn, raises philosophical questions about whether these models can be said to generate truly novel samples, or whether every output remains, in part, a rearrangement of prior data.
Figure 15: Examples of interpolations we observed, left: “Essential T-Shirt”, right: “iPhone Case & Cover”.
5.2 Perturbations
We also observed clusters of images that were nearly identical in their underlying template, yet differed by the manifestation of one or more objects with semantically similar alternatives placed in the same location—for example, swapping a lamp for a slightly different lamp, or a chair for another chair. An illustration of this behavior appears in Figure˜16. These perturbations preserve high CLIP similarity outside the editable region, while introducing small pixel-level differences that modestly reduce the $\ell_{2}$ similarity—though the images still remain far more similar than random pairs.
The precise origin of this phenomenon is difficult to isolate, but it aligns with two plausible explanations. First, minor perturbations to the latent noise can produce small, localized variations in the final image, consistent with behavior demonstrated in [17]. Second, it may reflect the one-step denoising behavior described by [37], which we observe specifically in the context of memorized images: once the global layout is recovered in the first step, subsequent steps allow only limited variation, resulting in small but consistent perturbations across generations.
5.3 Leakage
Another phenomenon we identified is what we term template leakage: cases where an image template belonging to one template group unexpectedly appears in generations prompted from another template group—or even from a prompt not associated with any template group at all. In these leaked generations, the reused template typically overlaps with the edited region in a visually coherent way. An illustrative example is shown in Figure˜17, where a template strongly associated with the “T-Shirt” template group appeared in a generation prompted for a “Tank Top.” Although the edited region correctly depicts a tank top, the surrounding background matches that of the T-shirt template observed in the source and in other reconstructed images.
However, inspecting e-commerce websites reveals that some templates are genuinely reused across product categories. As a result, not every suspicious case necessarily represents true leakage. For instance, Figure˜17 shows a background shared by both “Canvas Wall Art Print” and “Wall Tapestry” because the corresponding image template appears with both categories in the training data.
To differentiate actual leakage from legitimate multi-category reuse, we generated images from our fine-tuned model using prompts that were semantically similar (“Tea Cup”) and semantically dissimilar (“T-Shirt”) to the target category “Coffee Mug.” These additional generations revealed clear evidence of genuine leakage. In particular, with prompts such as “skg Tea Cup,” all resulting images displayed backgrounds that were either interpolations of known coffee-mug templates or matched their color and texture structure. Notably, the generated tea cups consistently appeared in top-view, a perspective absent from the training data and from all coffee-mug generations (which only exhibited the three canonical views). The cups also appeared in shapes not present in the train set. Taken together, these observations confirm that template leakage does occur, producing outputs that combine template structure from one template group with object semantics from another.
Background shared between template groups, not suspected for leakage as the source has been found for both. Left prompt: “Galaxy Wall Tapestry”, right prompt: “Galaxy Canvas Wall Art Print”.
Suspected leakage from “T-Shirt” to “Tank Top”. Left prompt: “Abstract Art Essential T-Shirt”, a source has been found. Right prompt: “Abstract Art Tank Top”, a source has not been found.
Figure 17: Example of a template shared between image categories in the training data, and suspected leakage between image categories.
Acknowledgments
Appendix A Identified Template Groups
Here we list the template groups that we identified through our method described on 3.
For convenience, the collocations are grouped by product category and the categories are ordered alphabetically.
The full list of collocations we identified as leading to the reconstruction of partially memorized visual content, is available on our git repository as a .txt file for convenient use.